OpenStudy (anonymous):

help please The table shows the outputs y for different inputs x: Input (x) 5 6 7 8 Output (y) 2 5 8 11 Part A: Do the data in this table represent a function? Justify your answer. Part B: Compare the data in the table with the relation f(x) = 2x + 13. Which relation has a greater value when x = 7? Part C: Using the relation in Part B, what is the value of x if f(x) = 75?

OpenStudy (unklerhaukus):

if each x, maps to only one y value, it can be a function

OpenStudy (unklerhaukus):

eg Input (x) 1 1 2 3 Output (y) 2 5 8 11 is not a function, because x = 1, gives two y values: 2 and 5

OpenStudy (unklerhaukus):

@Mr.kole does your table, have and x input, giving multiple y values?

OpenStudy (anonymous):

yh

OpenStudy (unklerhaukus):

pardon?

OpenStudy (unklerhaukus):

does any x input, give multiple y values?

OpenStudy (anonymous):

no

OpenStudy (unklerhaukus):

so can it be a function ?

OpenStudy (anonymous):

yh

OpenStudy (unklerhaukus):

yes.

OpenStudy (unklerhaukus):

now B What is the value in the table, that relates to x = 7

OpenStudy (anonymous):

hey i not reall good with algebra so please have mercy

OpenStudy (anonymous):

really

OpenStudy (unklerhaukus):

nope, what is y, when x is 7 use the table

OpenStudy (anonymous):

8

OpenStudy (unklerhaukus):

x=7 Input | (x) 5 6 7 8 Output | (y) 2 5 8 11 | y=8 yes!

OpenStudy (unklerhaukus):

now what is f(x) = 2x + 13. when x = 7 ie f(7) = 2(7) + 13. = =

OpenStudy (unklerhaukus):

2 times 7 plus 13 equals ?

OpenStudy (anonymous):

noooo

OpenStudy (unklerhaukus):

f(7) = 2(7) + 13. = 14 +13 =

OpenStudy (anonymous):

27

OpenStudy (unklerhaukus):

So Which relation has a greater value when x = 7?

OpenStudy (anonymous):

y

OpenStudy (unklerhaukus):

when x= 7: y = 8 and f = 27 which is greater

OpenStudy (anonymous):

f

OpenStudy (unklerhaukus):

it's f

OpenStudy (unklerhaukus):

right?

OpenStudy (unklerhaukus):

now for C

OpenStudy (unklerhaukus):

it asks, for which x, is f(x) = 2x + 13 = 75

OpenStudy (unklerhaukus):

so solve 2x+13 = 75 for x

OpenStudy (anonymous):

yh

OpenStudy (anonymous):

31

OpenStudy (unklerhaukus):

done!

OpenStudy (anonymous):

thanks